A +100μC charged is place at the origin. A +200μC charge is place 2m above the origin and a -300μC charged is placed on 3m to the right of the origin. What is the magnitude and direction of the net force acting on the charge placed at the origin?



1
Expert's answer
2021-09-10T12:18:41-0400

"F_1 = \frac{9 \times 10^9 \times 100 \times 300 \times 10^{-12}}{9} = 30 \;N \\

F_2 = \frac{9 \times 10^9 \times 100 \times 200 \times 10^{-12}}{4} = 45 \;N \\

F = \sqrt{F_1^2 +F^2_2} = 54.08 \;N"


"θ=180 +α \\

= 180 + tan^{-1}(\frac{45}{30}) \\

= 236.31"